# Search for Long-lived Particles at Future Lepton Colliders Using Deep Learning Techniques

### v5

@@ 3.3 Graph Neural Network
In the GNN approach [46], low-level calorimeter and tracker hits from an event are clustered and then integrated into a heterogeneous event graph, encompassing both calorimeter-type and tracker-type nodes. We implement a simple clustering method based on spatial proximity, grouping hits around the most energetic ones within a predefined region. This method is intentionally chosen for its computational simplicity, aiming to quickly and effectively reduce data dimensionality without requiring extensive event reconstruction. Given that lepton collider environments are relatively clean compared to hadron colliders, more complex clustering algorithms offer limited improvements in performance but at increased computational cost. Sensitivity studies confirmed that varying clustering parameters, such as minimum-hit thresholds and cluster radii, had negligible impact on signal efficiency and background rejection efficiency, thereby validating the robustness and suitability of our approach. The clustering process helps to minimize the complexity of the graph and significantly reduces memory usage during training. Furthermore, nodes of the same detector type are comprehensively interconnected, facilitating the formation of edges within the graph. This structure optimizes the analysis by ensuring efficient data processing and interaction modeling between different components of the detector hits. For the clustering of hits in the calorimeter, a point cloud composition method based on energy screening and distance thresholding is introduced. The most energetic calorimeter hit is first identified as the core hit. Subsequently, hits within a predefined distance threshold of $50~{}mm$ from the core hit are merged, constructing a graphical node structure. It is required that the number of merged hits exceeds 3 and the total number of hits in a cluster exceeds 10. This procedure is then iterated for all hits in an event to construct the calorimeter-type nodes. Hits that fail to meet these requirements are not clustered and are excluded from the GNN. The momentum $p$ of each calorimeter hit is defined as parallel to its position: $p_{i}=\frac{i}{r}E$, where $i$ represents the position along the $x$, $y$, or $z$ axis, with $z$ as the beam line direction, $x$ and $y$ in the transverse plane, and $y$ pointing towards the ground. $r$ is the radial distance $r=\sqrt{x^{2}+y^{2}+z^{2}}$, and $E$ is the hit’s energy. The tracker hits are binned into $5~{}\times~{}6$ blocks based on their R-$\phi$ positions. In each block, the tracker hits are clustered into a tracker-type node taking an arithmetic average of all hit positions. The definitions of nodes and edges for the GNN event graph are summarized in Table 3.
@@
Event graphs are then input into a GNN-based heterogeneous architecture, as illustrated in Figure 5. Features of different types of nodes and edges are first embedded into a high-dimension latent space and then forwarded to a Heterogeneous Detector Information Block (HDIB). The HDIB consists of two Detector Information Blocks (DIBs) [47, 48]and two multilayer perceptrons (MLPs). The HDIB adopts a parameter-sharing design between the tracker and the calorimeter types, specifically tailored for integrating information from different types of detectors. After the total $L$ layers of HDIBs, the final node embeddings of the tracker and the calorimeter are aggregated and forwarded to the decoding layer to form the classification scores.
@@ 4 Systematic Uncertainties
The dominant uncertainty source comes from the training variability of neural networks. To assess the robustness of our machine-learning models, 50 CNNs are trained with different initial seeds. The training uncertainty for the neural networks is estimated as half of the difference between the maximum and minimum efficiencies observed, which amounts to approximately $1.7\%$ 11 A similar procedure was applied to the GNN models, yielding a training uncertainty of approximately 2.3%, although this is not included in the final results since the CNN model serves as the baseline model.
. The uncertainty originating from the integrated luminosity for CEPC at $\sqrt{s}=240$ GeV is estimated to be $0.13\%$ [49], which is negligible in this study.
@@ 5 Results
The 95% confidence level upper limits on the branching ratio $\mathcal{B}(H\rightarrow XX)$ are derived using the CLs method [50], comparing the observed test statistics against the expected distributions under the null and alternative (with LLP signal) hypotheses. The analyzed samples correspond to a luminosity of $20~{}\text{ab}^{-1}$, corresponding to approximately $4\times 10^{6}$ Higgs bosons [51]. We consider two LLP signal scenarios:
@@
Table 5: The 95% C.L. exclusion limit on BR($h\to XX$) for all signal channels with both fixed and floating $\epsilon_{V}$. The limits include ±$\pm 1\sigma$ uncertainties after taking into account both statistical and systematic contributions.
Scenario $\mathcal{B}$ (×$\times 10^{-6}$) Lifetime [ns]
@@ 6.2 Comparisons with other searches
We evaluate the effectiveness of our ML-based approach for LLPs detection at future lepton colliders by comparing our results with results at hadron colliders with the ATLAS experiment [15, 14] and CMS experiment [16, 17] as well as future HL-LHC experiments [52]. The comparison is based on four primary metrics: signal acceptance, selection efficiency, analysis strategy and signal yields:
  • Signal Acceptance: Both ATLAS and CMS results have limited signal acceptance, typically a few percent, as they focus on LLPs decaying in the muon detector. In contrast, our ML-based approach covers the entire detector, resulting in 100% signal acceptance except for LLPs with long lifetime ($>10$ ns) and low mass ($<10$ GeV).
  • Selection Efficiency: For hadron colliders, LLPs events typically trigger on displaced decays and/or large missing transverse energy. The LLPs trigger efficiency at the ATLAS experiment is estimated to be between $10^{-3}$ and $0.3$ [15]. Besides trigger efficiency, there are additional efficiencies involved such as displaced vertex/object reconstruction efficiencies which are typically in the order of a few percent. In contrast, LLPs event selection at lepton colliders can adapt to a triggerless-equivalent approach [53, 54] owing to the clean environment. The ML-based approach can be applied directly with low-level detector information without any event-level reconstruction. As a result, our ML-based approach with lepton colliders can achieve an overall selection efficiency as high as $95\%$, an improvement of several orders in magnitude when compared with LHC or HL-LHC efficiencies.
@@
  • Signal Yields and Upper Limits Comparison: LHC and HL-LHC can produce a significantly larger number of Higgs bosons compared to lepton colliders. Despite this, higher signal acceptance and selection efficiencies in our ML-based approach compensate for the relatively low number of Higgs bosons. We achieve upper limits as low as $9.7\times 10^{-7}$ on $\mathcal{B}(H\rightarrow XX)$ with $4\times 10^{6}$ Higgs bosons. This upper limit is approximately three orders of magnitude better than the $10^{-3}$ limit observed at the ATLAS and CMS experiments with $10^{7}$ Higgs bosons and it is comparable to the projected HL-LHC limit with about $10^{8}$ expected Higgs bosons.
Besides comparing with LLPs searches at hadron colliders, we have also compared our result with a preliminary LLPs study [55] on the ILC [28] sensitivity with a traditional selection-based method. The ILC sensitivity study searches for long-lived dark photons produced in Higgstrahlung events via the Higgs portal. We have compared our result with the hadronic decay dark photo result since the event signature is similar. We have seen that the signal acceptance factors are similar between ILC and CEPC detectors but the signal efficiencies differ significantly. The signal efficiencies in the ILC study range from $0.1\%$ to $10\%$, which is at least an order of magnitude lower than ours. The upper limits on $\mathcal{B}(H\rightarrow XX)$ in the ILC study are derived under the assumption of $100\%$ truth-level signal efficiency. Under this assumption, two results show similar sensitivities only in the low lifetime region ($<1~{}ns$) of LLPs. In the long lifetime region, for example, for a 1 GeV LLP with a lifetime of a few nanoseconds, our result yields an upper limit of about $5\times 10^{-6}$ which is an order of magnitude better than the ILC’s upper limit of $10^{-4}$.
@@ 7 Conclusion
In summary, we have investigated the search for LLPs at future lepton colliders using machine learning techniques with full simulation data samples. We have demonstrated that the efficiency of LLP signal detection can reach up to 95% for LLPs with a mass around 50 GeV and a lifetime of approximately 1 nanosecond, surpassing traditional selection-based methods on almost all fronts. Our methodology’s effectiveness is not confined to theoretical constructs but is evidenced by detailed simulations. With a luminosity of $20~{}\text{ab}^{-1}$ and around $4\times 10^{6}$ Higgs bosons analyzed, this study sets new benchmarks for sensitivity in LLP searches with lepton colliders, the branching ratio of Higgs decaying into LLPs reaches $9.7\times 10^{-7}$.
@@
Acknowledgement The authors thank theorists Zhen Liu, Tao Liu, Jia Liu, Ziping Wang, as well as experimentalists Manqi Ruan and Gang Li for their useful suggestions and inputs about the paper. This work was supported by National Key R&D Program of China (Grant No. 2024YFA1610603, 2023YFA1606003); National Natural Science Foundation of China (Grants No. 12305217, 12375072 and 12475108).
@@ Appendix B Pile-up and cosmic ray backgrounds
Given the parameters for the CEPC [42], with $L=5\times 10^{34}\,\textrm{cm}^{-2}\textrm{s}^{-1}$, $\sigma=1000\times 10^{-36}\,\textrm{cm}^{2}$, and $\tau=636\times 10^{-9}\,\textrm{s}$, the computed pile-up rate is approximately $3\times 10^{-5}$. This low pile-up rate underscores the minimal impact of pile-up on the CEPC’s operation, making the pile-up background negligible.
The cosmic ray background, particularly from muons, is effectively mitigated in lepton colliders through several ways, making it a negligible concern in most analyses. Firstly, the timing method uses the fact that cosmic muons do not coincide with the precise timing of collision events, allowing for a clear differentiation based on the synchronization of particle detection with beam crossings. Secondly, the distinct trajectories of cosmic muons, which are typically vertical, are different from the varied directions of muons originating from collision points. Moreover, the underground placement of detectors significantly shields them from cosmic interference, complemented by additional physical barriers that absorb cosmic radiation. The combination of these strategies effectively minimizes the cosmic ray background noise. This background is also ignored in other LLPs search experiments such as the MATHUSLA experiment [56].
@@ Appendix F Sensitivity with different background estimation
Table 6: The 95% C.L. exclusion limit on BR($h\to XX$) for all signal channels with both fixed and floating $\epsilon_{V}$ with 2 backgrounds assumption. The limits include ±$\pm 1\sigma$ uncertainties after taking into account both statistical and systematic contributions.
Scenario $\mathcal{B}$ (×$\times 10^{-6}$) Lifetime [ns]
@@ Appendix G Comparison with selection-based analyses
[…]
@@
Traditional methods generally faces various difficulties in reconstructing and identifying both Z bosons and LLPs, and thus require complicated case-by-case analyses, while the machine learning approach does not. The latter approach uniformly makes use of all available information of event data, and not only exhibits a technical superiority but also systematically improves signal efficiencies. Comparing Table 4 and Table 7, the ML-based signal efficiency is about 97%-98% with a LLP lifetime of 10 ns and mass of 50 GeV, significantly higher than 50.8% obtained with the selection-based approach.
We have also compared our results with a previous LLP search [57] conducted at lepton colliders, where the selection-based method is implemented with fine-tuning and optimization for varying scenarios, achieving a maximum selection efficiency ***”Selection efficiency” specifically denotes the product of signal acceptance and signal efficiency. of 0.76 for a 50 GeV particle with a 1 ns lifetime. In contrast, our ML approach significantly outperforms this result and achieves a maximum efficiency of 0.99. Notably, our method maintained high selection efficiencies (over 0.9) for lifetimes under 1 ns and masses above 10 GeV, and reasonable efficiencies (around 0.4) for masses as low as 1 GeV. Furthermore, while the selection-based analysis is limited to lifetimes of up to 1 ns, our ML approach can reach up to 100 ns, demonstrating its superior capability in performing a broader spectrum of LLP searches.
@@
Selections LLPs Signal with $Z\to j\bar{j}$ $ee\to q\bar{q}$ $ee\to ZH$

### v6

@@ 3.3 Graph Neural Network
In the GNN approach [Shlomi_2021], low-level calorimeter and tracker hits from an event are clustered and then integrated into a heterogeneous event graph, encompassing both calorimeter-type and tracker-type nodes. We implement a simple clustering method based on spatial proximity, grouping hits around the most energetic ones within a predefined region. This method is intentionally chosen for its computational simplicity, aiming to quickly and effectively reduce data dimensionality without requiring extensive event reconstruction. Given that lepton collider environments are relatively clean compared to hadron colliders, more complex clustering algorithms offer limited improvements in performance but at increased computational cost. Sensitivity studies confirmed that varying clustering parameters, such as minimum-hit thresholds and cluster radii, had negligible impact on signal efficiency and background rejection efficiency, thereby validating the robustness and suitability of our approach. The clustering process helps to minimize the complexity of the graph and significantly reduces memory usage during training. Furthermore, nodes of the same detector type are comprehensively interconnected, facilitating the formation of edges within the graph. This structure optimizes the analysis by ensuring efficient data processing and interaction modeling between different components of the detector hits. For the clustering of hits in the calorimeter, a point cloud composition method based on energy screening and distance thresholding is introduced. The most energetic calorimeter hit is first identified as the core hit. Subsequently, hits within a predefined distance threshold of 50mm50~$m$ from the core hit are merged, constructing a graphical node structure. It is required that the number of merged hits exceeds 3 and the total number of hits in a cluster exceeds 10. This procedure is then iterated for all hits in an event to construct the calorimeter-type nodes. Hits that fail to meet these requirements are not clustered and are excluded from the GNN. The momentum $p$ of each calorimeter hit is defined as parallel to its position: $p_{i}=\frac{i}{r}E$, where $i$ represents the position along the $x$, $y$, or $z$ axis, with $z$ as the beam line direction, $x$ and $y$ in the transverse plane, and $y$ pointing towards the ground. $r$ is the radial distance $r=\sqrt{x^{2}+y^{2}+z^{2}}$, and $E$ is the hit’s energy. The tracker hits are binned into $5~\times~6$ blocks based on their R-$\phi$ positions. In each block, the tracker hits are clustered into a tracker-type node taking an arithmetic average of all hit positions. The definitions of nodes and edges for the GNN event graph are summarized in Table 3.
@@
Event graphs are then input into a GNN-based heterogeneous architecture, as illustrated in Figure 5. Features of different types of nodes and edges are first embedded into a high-dimension latent space and then forwarded to a Heterogeneous Detector Information Block (HDIB). The HDIB consists of two Detector Information Blocks (DIBs) [lorentznet, satorras2021n]and two multilayer perceptrons (MLPs). The HDIB adopts a parameter-sharing design between the tracker and the calorimeter types, specifically tailored for integrating information from different types of detectors. After the total $L$ layers of HDIBs, the final node embeddings of the tracker and the calorimeter are aggregated and forwarded to the decoding layer to form the classification scores.
@@ 4 Systematic Uncertainties
The dominant uncertainty source comes from the training variability of neural networks. To assess the robustness of our machine-learning models, 50 CNNs are trained with different initial seeds. The training uncertainty for the neural networks is estimated as half of the difference between the maximum and minimum efficiencies observed, which amounts to approximately $7\% 11$ A similar procedure was applied to the GNN models, yielding a training uncertainty of approximately 2.3%, although this is not included in the final results since the CNN model serves as the baseline model.
.The theoretical uncertainty of the ZH cross section is estimated to be 3% [Freitas:2022hyp]. The uncertainty originating from the integrated luminosity for CEPC at $\sqrt{s}=240$ GeV is estimated to be $0.13\%$ [lum_sys], which is negligible in this study.
@@ 5 Results
The 95% confidence level upper limits on the branching ratio $\mathcal{B}(H\rightarrow XX)$ are derived using the CLs method [CLs], comparing the observed test statistics against the expected distributions under the null and alternative (with LLP signal) hypotheses. The analyzed samples correspond to a luminosity of $20~\text{ab}^{-1}$, corresponding to approximately $4\times 10^{6}$ Higgs bosons [CEPCStudyGroup:TDR1]. We consider two LLP signal scenarios:
@@
Table 5: The 95% C.L. exclusion limit on BR($h\to XX$) for all signal channels with both fixed and floating $\epsilon_{V}$. The limits include $\pm 1\sigma$ uncertainties after taking into account both statistical and systematic contributions.
Scenario $\mathcal{B}$ ($\times 10^{-6}$) Lifetime [ns]
@@ 6.2 Comparisons with other searches
We evaluate the effectiveness of our ML-based approach for LLPs detection at future lepton colliders by comparing our results with results at hadron colliders with the ATLAS experiment [ATLAS2024LLP, ATLAS_llp_muon] and CMS experiment [CMS:llp2023, CMS_llp_2024] as well as future HL-LHC experiments [HL-LHC]. The comparison is based on four primary metrics: signal acceptance, selection efficiency, analysis strategy and signal yields:
  • Signal Acceptance: Both ATLAS and CMS results have limited signal acceptance, typically a few percent, as they focus on LLPs decaying in the muon detector. In contrast, our ML-based approach covers the entire detector, resulting in 100% signal acceptance except for LLPs with long lifetime (>10>10 ns) and low mass (<10<10 GeV).
  • Selection Efficiency: For hadron colliders, LLPs events typically trigger on displaced decays and/or large missing transverse energy. The LLPs trigger efficiency at the ATLAS experiment is estimated to be between $10^{-3}$ and $0.3$ [ATLAS2024LLP]. Besides trigger efficiency, there are additional efficiencies involved such as displaced vertex/object reconstruction efficiencies which are typically in the order of a few percent. In contrast, LLPs event selection at lepton colliders can adapt to a triggerless-equivalent approach [trigger1, trigger2] owing to the clean environment. The ML-based approach can be applied directly with low-level detector information without any event-level reconstruction. As a result, our ML-based approach with lepton colliders can achieve an overall selection efficiency as high as $95\%$, an improvement of several orders in magnitude when compared with LHC or HL-LHC efficiencies.
@@
  • Signal Yields and Upper Limits Comparison: LHC and HL-LHC can produce a significantly larger number of Higgs bosons compared to lepton colliders. Despite this, higher signal acceptance and selection efficiencies in our ML-based approach compensate for the relatively low number of Higgs bosons. We achieve upper limits as low as $1.0\times 10^{-6}$ on $\mathcal{B}(H\rightarrow XX)$ with $4\times 10^{6}$ Higgs bosons. This upper limit is approximately three orders of magnitude better than the $10^{-3}$ limit observed at the ATLAS and CMS experiments with $10^{7}$ Higgs bosons and it is comparable to the projected HL-LHC limit with about $10^{8}$ expected Higgs bosons.
Besides comparing with LLPs searches at hadron colliders, we have also compared our result with a preliminary LLPs study [jeanty2022sensitivity] on the ILC [ILC:2019gyn] sensitivity with a traditional selection-based method. The ILC sensitivity study searches for long-lived dark photons produced in Higgstrahlung events via the Higgs portal. We have compared our result with the hadronic decay dark photo result since the event signature is similar. We have seen that the signal acceptance factors are similar between ILC and CEPC detectors but the signal efficiencies differ significantly. The signal efficiencies in the ILC study range from $0.1\%$ to $10\%$, which is at least an order of magnitude lower than ours. The upper limits on $\mathcal{B}(H\rightarrow XX)$ in the ILC study are derived under the assumption of $100\%$ truth-level signal efficiency. Under this assumption, two results show similar sensitivities only in the low lifetime region (<1ns<1~ns) of LLPs. In the long lifetime region, for example, for a 1 GeV LLP with a lifetime of a few nanoseconds, our result yields an upper limit of about $5\times 10^{-6}$ which is an order of magnitude better than the ILC’s upper limit of $10^{-4}$.
@@ 7 Conclusion
In summary, we have investigated the search for LLPs at future lepton colliders using machine learning techniques with full simulation data samples. We have demonstrated that the efficiency of LLP signal detection can reach up to 95% for LLPs with a mass around 50 GeV and a lifetime of approximately 1 nanosecond, surpassing traditional selection-based methods on almost all fronts. Our methodology’s effectiveness is not confined to theoretical constructs but is evidenced by detailed simulations. With a luminosity of $20~\text{ab}^{-1}$ and around $4\times 10^{6}$ Higgs bosons analyzed, this study sets new benchmarks for sensitivity in LLP searches with lepton colliders, the branching ratio of Higgs decaying into LLPs reaches $1.0\times 10^{-6}$.
@@
8 Acknowledgement
The authors thank theorists Zhen Liu, Tao Liu, Jia Liu, Ziping Wang, as well as experimentalists Manqi Ruan and Gang Li for their useful suggestions and inputs about the paper. This work was supported by National Key R&D Program of China (Grant No. 2024YFA1610603, 2023YFA1606003); National Natural Science Foundation of China (Grants No. 12305217, 12375072 and 12475108).
@@ Appendix B Pile-up and cosmic ray backgrounds
Given the parameters for the CEPC [CEPCStudyGroup:CDR2], with $L=5\times 10^{34}\,\textrm{cm}^{-2}\textrm{s}^{-1}$, $\sigma=1000\times 10^{-36}\,\textrm{cm}^{2}$, and $\tau=636\times 10^{-9}\,\textrm{s}$, the computed pile-up rate is approximately $3\times 10^{-5}$. This low pile-up rate underscores the minimal impact of pile-up on the CEPC’s operation, making the pile-up background negligible.
The cosmic ray background, particularly from muons, is effectively mitigated in lepton colliders through several ways, making it a negligible concern in most analyses. Firstly, the timing method uses the fact that cosmic muons do not coincide with the precise timing of collision events, allowing for a clear differentiation based on the synchronization of particle detection with beam crossings. Secondly, the distinct trajectories of cosmic muons, which are typically vertical, are different from the varied directions of muons originating from collision points. Moreover, the underground placement of detectors significantly shields them from cosmic interference, complemented by additional physical barriers that absorb cosmic radiation. The combination of these strategies effectively minimizes the cosmic ray background noise. This background is also ignored in other LLPs search experiments such as the MATHUSLA experiment [MATHUSLA_2019].
@@ Appendix F Sensitivity with different background estimation
Table 6: The 95% C.L. exclusion limit on BR($h\to XX$) for all signal channels with both fixed and floating $\epsilon_{V}$ with 2 backgrounds assumption. The limits include $\pm 1\sigma$ uncertainties after taking into account both statistical and systematic contributions.
Scenario $\mathcal{B}$ ($\times 10^{-6}$) Lifetime [ns]
@@ Appendix G Comparison with selection-based analyses
  • $\not${E}: the missing energy determined with reconstructed energy in detector subtraction from initial electron-positron c.m. energy.
@@
Traditional methods generally faces various difficulties in reconstructing and identifying both Z bosons and LLPs, and thus require complicated case-by-case analyses, while the machine learning approach does not. The latter approach uniformly makes use of all available information of event data, and not only exhibits a technical superiority but also systematically improves signal efficiencies. Comparing Table 4 and Table 7, the ML-based signal efficiency is about 91%-95% with a LLP lifetime of 10 ns and mass of 50 GeV, significantly higher than 50.8% obtained with the selection-based approach.
We have also compared our results with a previous LLP search [LLP_2019] conducted at lepton colliders, where the selection-based method is implemented with fine-tuning and optimization for varying scenarios, achieving a maximum selection efficiency ***”Selection efficiency” specifically denotes the product of signal acceptance and signal efficiency. of 76% for a 50 GeV particle with a 1 ns lifetime. In contrast, our ML approach significantly outperforms this result and achieves a maximum efficiency of 95%. Notably, our method maintained high selection efficiencies (over 0.9) for lifetimes under 1 ns and masses above 10 GeV, and reasonable efficiencies (around 0.4) for masses as low as 1 GeV. Furthermore, while the selection-based analysis is limited to lifetimes of up to 1 ns, our ML approach can reach up to 100 ns, demonstrating its superior capability in performing a broader spectrum of LLP searches.
@@
Selections LLPs Signal with $Z\to j\bar{j}$ $e\to q\bar{q} ee$$e\to ZH$